AlexanderV:

The line segment on the map below is 1.6 cm long, and the map scale is 1 cm = 0.5 km. What is the gradient of the slope marked by the line?

AlexanderV:

H. E. L. P.

jhonyy9:

so than you know the rule of 3 simple you can use it and write 1 cm mean 0,5 km 1,6 cm mean x km so cross multiplie and calcule the x value x = 1,6 *0,5 /1 = 1,6 *0,5 = ? km hope helped understandably easy

AlexanderV:

0.8